OpenStudy (anonymous):

always, sometimes, or never? x + x = 2x

OpenStudy (anonymous):

always...

OpenStudy (anonymous):

always.

OpenStudy (anonymous):

x + x = 2*x = 2x.

OpenStudy (saifoo.khan):

Always man!